slipknot3066:

I got 15 questions in Algebra I could use any help I could get

slipknot3066:

1 attachment
dude:

First divide the coefficients, then you subtract the value of the expoenents Rule when dividing exponents (They must have the same base) \(\large\frac{d^{7}}{d^{3}}= d^{7-3}=d^4\)